# Two-Photon Absorption in Silicon Using Real Density Matrix Approach

### v1

@@ Abstract
Two-photon absorption in indirect gap semiconductors is an frequently encountered, but not well-understood phenomenon. To address this, the Real Density Matrix Approach is applied to describe two-photon absorption in silicon through the excitonic response to the interacting fields. This approach produces an analytical expression for the dispersion of the two-photon absorption coefficient for indirect-gap materials, and can be used to explain trends in reported experimental data for bulk silicon both old and new with minimal fitting.
preprint: APS/123-QED
@@ II.1 Real density matrix approach
In the RDMA approach the nonlinear response is be described by a set of coupled equations for the exciton density represented by the coherent amplitude $Y({\bf r}_{1},{\bf r}_{2})$, the density matrix for the conduction electrons $C({\bf r}_{1},{\bf r}_{2})$, and the density matrix $D({\bf r}_{1},{\bf r}_{2})$ for the holes in the valence band. [24]. These matrices are sub-matrices of the following density matrix
$\underline{\underline{\hat{\rho}}}=\begin{pmatrix}C&Y^{*}\\ Y&1-D\end{pmatrix}.$ (3)
The evolution of the system is described by Heisenberg equation for the density matrix $\frac{\delta\underline{\underline{\hat{\rho}}}}{\delta t}=[\underline{\underline{\hat{\rho}}},\hat{H}]$, supplemented by relaxation terms. We also define the excitonic center-of-mass coordinate
${\bf R}={\bf R}_{12}=\frac{m_{h}{\bf r}_{1}+m_{e}{\bf r}_{2}}{m_{h}+m_{e}},$ (4)
and ${\bf r}={\bf r}_{1}-{\bf r}_{2}$ is the electron-hole relative coordinate.
The Hamiltonian consists of two parts
$\hat{H}=H_{eh}+H^{\prime},$ (5)
where $H_{eh}$ is the electron-hole Hamiltonian [24], and $H^{\prime}$ describes the exciton-phonon interaction
@@
$V_{p}({\bf q},{\bf r})=\left(\frac{\hbar}{2M\omega_{ph}}\right)^{1/2}$
×$\times\sum\limits_{{\bf R}_{a}}e^{i\textbf{q}\textbf{R}_{a}}\textbf{e}_{0}\hbox{\boldmath$\nabla$}_{r}V({\bf r}-{\bf R}_{a}).$ (7)
The $a({\bf q})$ is the phonon annihilation operator, $V({\bf r}-{\bf R}_{a})$ is the potential at the point ${\bf R_{a}}$, $e_{0}$ is the phonon polarization, $\omega_{ph}$ is the phonon frequency and q is its wave vector.
The constitutive equations for the $Y,C$ and $D$ matrices have the following form
${i}\hbar\partial_{t}Y-H\,Y=-{\bf M}{\bf E}$
$+{\bf E}{\bf M}_{0}C+{\bf E}{\bf M}_{0}D+{i}\hbar\left(\frac{\partial Y}{\partial t}\right)_{{\rm irrev}},$ (8)
${i}\hbar\partial_{t}C+H_{ee}C={\bf M}_{0}{\bf E}(Y-Y^{*})+{i}\hbar\left(\frac{\partial C}{\partial t}\right)_{{\rm irrev}},$ (9)
${i}\hbar\partial_{t}D-H_{hh}D={\bf M}_{0}{\bf E}(Y-Y^{*})+{i}\hbar\left(\frac{\partial D}{\partial t}\right)_{{\rm irrev}},$ (10)
where the smeared-out transition dipole density ${\bf M}({\bf r})$ and its integrated strength ${\bf M}_{0}$ are related to the bilocality of the amplitude $Y$ and describe the quantum coherence between the macroscopic electromagnetic field and the inter-band transitions. The electron-electron and hole-hole interaction Hamiltonians $H_{ee}$ and $H_{hh}$ were defined in reference [24]. In expressions (8)-(10), the electromagnetic field $\bf E$ includes two frequencies $\omega_{a}$ and $\omega_{b}$, and is written as
${\bf E}=\textbf{E}_{0a}\exp(i\textbf{k}_{\textbf{a}}{\textbf{R}}-i\omega_{a}t)$
[…]
The terms $(\frac{\partial...}{\partial t})_{{\rm irrev}}$ describe the dissipation and relaxation processes, and satisfy
$\left.\frac{\partial\underline{\underline{\hat{\rho}}}}{\partial t}\right|_{irrev}=$ (12)
$\quad=-\begin{pmatrix}\frac{1}{T_{1}}[C(t)-C^{(0)}]&\frac{1}{T_{2}}[Y^{*}(t)-Y^{*(0)}]\cr\frac{1}{T_{2}}[Y(t)-Y^{(0)}]&\frac{1}{T_{1}}[D(t)-D^{(0)}]\end{pmatrix}.$ (15)
where the states with superscript $(0)$ represent the steady state solutions. The times $T_{1},T_{2}$ describe the characteristic duraction of all dephasing processes: electron-electron, electron-phonon, and electron-photon interactions. Here, $T_{1},T_{2}$ are taken as phenomenological constants.
The excitonic part of the medium’s polarization is obtained from
${\bf P}({\bf R},t)=2\int{d}^{3}r\,\textbf{M}^{*}({\bf r})\hbox{Re}~{}Y({\bf R},{\bf r},t)$
[…]
The linear optical properties are calculated by solving the interband equation (8), supplemented by the corresponding Maxwell equation, where the polarization (II.1) acts as a source. For computing the nonlinear optical properties we use the entire set of constitutive equations (8)-(10). Although finding a general solution of the equations is challenging, in special situations a solution can be found. For example, if one assumes that the matrices $Y,C$ and $D$ can be expanded in powers of the electric field ${\bf E}$, an iterative procedure can be used.
@@
    In the energy range $\hbar\omega_{a}\;+\hbar\omega_{b}\,>2E_{g}$, we solve the equations (19), (40) and following equations for the matrices $C,D$ assuming $V_{eh}=0$ [24], thus entering the energy range represented by continuum states. The solution is obtained in terms of an appropriate Green function.
II.2 Discrete states
Our goal is to derive expressions for the NTA absorption coefficients. These can be obtained from the third-order nonlinear susceptibility, which, in turn, can be determined via an iterative procedure within the context of the RDMA. The first step in the iteration consists of solving the equation (8), which at this stage takes on the form

### v2

# Two-Photon Absorption in Silicon Using the Real Density Matrix Approach

@@ Abstract
Two-photon absorption in indirect gap semiconductors is an frequently encountered, but not well-understood phenomenon. To address this, the real-density matrix approach (RDMA) is applied to describe two-photon absorption in silicon through the excitonic response to the interacting fields. This approach produces an analytical expression for the dispersion of the two-photon absorption coefficient for indirect-gap materials, and can be used to explain trends in reported experimental data for bulk silicon both old and new with minimal fitting.
preprint: AIP/123-QED
@@ II.1 Real density matrix approach
In this section, we briefly review the basic principles of the RDMA method in the context of linear and nonlinear optical excitations in semiconductors. In general, the optical response of semiconductors to incoming electromagnetic (EM) waves can be described in terms of the correlations in a many-body system (semiconductor) caused by the interactions between the EM fields, electrons, and holes (quasiparticles). A variety of theoretical methods has been used to describe the response, including the Green function approach [26, 27] and the RDMA. The latter, also known as coherent wave theory or the band edge equations, was developed in works by Stahl, Balslev, Axt, Zimmermann and others, see for instance Refs. [28, 29].
In this work, we have chosen the RDMA to take advantage of its formulation in the real space. The method provides a direct relation between the density matrices and relevant observables, allowing an easy comparison between experimental and theoretical results. In general, the Coulomb interaction between the carriers in a many-body system produces an infinite hierarchy of evolution equations for $n-$point density matrices. The lowest level consists of two-point density matrices, which describe the interband transitions, between the valence and the conduction band, as well as the intraband transitions. Two-point density matrices are directly related to measurable quantities such as polarization and carrier densities, whereas higher order correlations are related, for example, to the formation of biexcitons. [30]
Below we restrict our model to two-point density matrices. The basic equations of the chosen approach are called the constitutive equations, or band-edge equations, and will be applied to describe the 2PA processes. We follow the description developed in Ref. [29]. The RDMA takes into account the following contributions: a) the electron-hole interaction; b) the dipole interaction between the electron-hole pairs and the electromagnetic field; c) the particle-surface interaction; and d) effects of external fields. We consider a semiconductor in the real space representation, characterized by a number of valence- and conduction bands. Electrons at site $j$ in the conduction band are described by fermion operators $\hat{c}^{c{\dagger}}_{j}(\hat{c}^{c}_{{j}})$, which correspond to the creation (annihilation) operators. Similarly, operators $\hat{d}^{v{\dagger}}_{{j}}(\hat{d}^{v}_{{j}})$ are creation (annihilation) operators for holes in valence bands at site $j$. In the case of direct interband transitions, the Hamilton operator in our model consists of three parts
$H=H_{0}+H_{em}+{H}_{C}.$ (1)
The term ${H}_{0}$ idnicates the one-particle Bloch states in the conduction and valence bands, and describes the intraband transport processes:
${H}_{0}=\sum_{j\ell}\left(\sum\limits_{c}T^{c}_{\ell j}\hat{c}_{\ell}^{c{\dagger}}\hat{c}_{j}^{c}-\sum\limits_{v}T^{v}_{j\ell}\hat{d}_{\ell}^{v{\dagger}}\hat{d}_{j}^{v}\right),$ (2)
where the transfer matrices $T_{ij}^{c,v}$ are given in terms of the band eigenvalues $E_{c,v}(k)$ as:
$T_{ij}^{c}=\frac{1}{N}\sum\limits_{k}e^{ik(R_{i}-R_{j})}E_{c}(k),$ (3)
with a similar expression for the valence band transfer matrix $T_{ij}^{v}$. The operator $H_{em}$ describes the interaction with the electromagnetic field:
$H_{em}=-\sum\limits_{j\ell}{\bf E}_{j\ell}\cdot\sum_{cv}{\bf M}^{cv*}_{j\ell}\hat{d}^{v}_{\ell}\hat{c}^{c}_{j}+{\rm h.c.},$ (4)
and ${\bf M}^{cv}_{j\ell}$ is the interband-dipole matrix element between Wannier states, which can be cast in the following form
${\bf M}^{cv}_{j\ell}=\frac{1}{N}\sum_{k}e^{ik(R_{j}-R_{\ell})}{\bf M}_{cv}({\bf k}),$ (5)
in terms of the dipole matrix element between Bloch functions
${\bf M}_{cv}({\bf k})=-e\langle\psi_{c}({\bf k})|{\bf r}|\psi_{v}({\bf k})\rangle.$ (6)
The electric field ${\bf E}_{j\ell}$ is considered at a mean point between the sites $j$ and $\ell$ of the lattice; below this position is assumed to coincide with the exciton center-of-mass. In the following, we assume the band structure is known from specific calculations.
The carrier interactions are described by the term ${H}_{C}$ of the Hamiltonian:
${H}_{C}=\frac{1}{2}\sum\limits_{i\neq j}V_{ij}(\hat{n}_{i}-\hat{h}_{i})(\hat{n}_{j}-\hat{h}_{j}),$ (7)
where the electron-hole interaction is screened by a background dielectric constant $\epsilon_{b}$
$V_{ij}=\frac{e^{2}}{4\pi\epsilon_{0}\epsilon_{b}\left|\textbf{r}_{i}-\textbf{r}_{j}\right|},$ (8)
We also define the electron and hole occupation numbers as:
$\hat{n}_{j}=\sum\limits_{c}\hat{c}_{j}^{c{\dagger}}\hat{c}_{j}^{c},\quad\hat{h}_{j}=\sum\limits_{v}\hat{d}^{v{\dagger}}_{j}\,\hat{d}^{v}_{j}.$ (9)
The physical quantities relevant for finding the optical properties can be expressed in terms of mean values of the following pair operators
excitonic transition density amplitude:
$Y_{12}^{\alpha\,b}=\langle\hat{Y}_{12}^{\alpha\,b}\rangle=\langle\hat{d}^{\alpha}_{1}\,\hat{c}^{b}_{2}\rangle,$
[…]
where the indices $a,b,..$ label the conduction bands while $\alpha,\beta,...$ label the valence bands. The excitonic transition density $Y^{\alpha\,b}_{12}$ contributes to the interband transition polarization with the following term
$P=2\,\hbox{Re}\,\left(\int_{r=r_{1}-r_{2}}d^{3}r\,\sum_{cv}\,M_{21}^{cv*}Y_{12}^{vc}\right),$ (11)
and the diagonal elements (the matrices $C$ and $D$) correspond to the densities of electrons
$\rho_{e}=\left.-e\sum_{c}\,C^{cc}_{12}\right|_{r_{1}=r_{2}},$ (12)
and holes
$\rho_{h}=\left.e\sum_{v}\,D^{vv}_{12}\right|_{r_{1}=r_{2}}.$ (13)
The matrices above are sub-matrices of the following density matrix
$\underline{\underline{\hat{\rho}}}=\begin{pmatrix}C_{cc^{\prime}}&Y^{*}_{vc}\cr Y_{vc}&1-D_{vv^{\prime}}\end{pmatrix}.$ (14)
The dynamics of the two-point matrices $Y,C,D$ is part of the hierarchy of reduced density matrices, and is obtained from the Heisenberg equations of motion
$i\hbar\partial_{t}\underline{\underline{\hat{\rho}}}=[\underline{\underline{\hat{\rho}}},{H}]+i\hbar\partial_{t}\underline{\underline{\dot{\hat{\rho}}}}_{\rm irrev},$ (15)
where the term $\underline{\underline{\dot{\hat{\rho}}}}_{\rm irrev}$ describes the irreversible dissipation and radiation decay processes due to all dephasing processes. In this paper, we consider electron-electron interactions, electron-phonon interactions and the optical transitions caused by the electromagnetic field. In many practical calculations all irreversible processes are described in terms of two dephasing times $T_{1},T_{2}$ which are taken as phenomenological constants that satisfy the following equation
$\left.\frac{\partial\underline{\underline{\hat{\rho}}}}{\partial t}\right|_{\rm{irrev}}=$ (16)
$\quad=-\begin{pmatrix}\frac{1}{T_{1}}[C(t)-C^{(0)}]&\frac{1}{T_{2}}[Y^{*}(t)-Y^{*(0)}]\cr\frac{1}{T_{2}}[Y(t)-Y^{(0)}]&\frac{1}{T_{1}}[D(t)-D^{(0)}]\end{pmatrix}.$
where the states with superscript ${(0)}$ denote the steady state solutions. The Hamiltonian in Eq. (15) can be written as
$H=H_{ee}+H_{hh}+H_{eh}+H_{ph}$ (17)
where we have the electron-hole interaction
$H_{eh}=E_{g}-V_{12}+\frac{1}{2m_{h}}\left(p_{1}-eA_{1}\right)^{2}+\frac{1}{2m_{e}}\left(p_{2}+eA_{2}\right)^{2}$
$+e\left(\Phi^{h}_{1}-\Phi^{e}_{2}\right),$ (18)
the electron Hamiltonian
$H_{ee}=\frac{1}{2m_{e}}\left[(p_{2}+eA_{2})^{2}-(p_{1}-eA_{1})^{2}\right]+e\left(\Phi^{e}_{1}-\Phi^{e}_{2}\right),$ (19)
and the hole Hamiltonian
$H_{hh}=\frac{1}{2m_{h}}\left[(p_{2}-eA_{2})^{2}-(p_{1}+eA_{1})^{2}\right]-e\left(\Phi^{h}_{1}-\Phi^{h}_{2}\right)$ (20)
where $E_{g}$ denotes the gap energy, $m_{e},m_{h}$ are the effective masses of electrons and holes, respectively, $-V_{12}$ is the statically screened Coulomb potential, $M_{0}$ is the element in equation (5) integrated over the real space, $A_{j}$ the vector potential of EM field at position ${\bf r}_{j}$, which may include an external magnetic field, $\Phi^{e/h}_{j}$ the scalar, external or electromagnetically induced potential acting on electrons (or holes) at position ${\bf r}_{j}$, and $E_{j}$ denotes the electric field of the radiation at the point ${\bf r}_{j}$. We neglect for the moment the vectorial and tensorial indices, and use the common notation for the momentum operators: $p_{1}=-i\hbar\nabla_{1}$ etc. The electron-electron, hole-hole and electron-hole exchange terms are included in the third term of the Hamiltonian (1), as given by expression (7), and are obtained by means of the random phase approximation (RPA) decoupling scheme. [31]
The exciton-phonon interaction is given by
[…]
with
$V_{p}({\bf q},{\bf r})=\sum\limits_{{\bf R}_{a}}\left(\frac{\hbar}{2M\omega_{ph}}\right)^{1/2}e^{i\textbf{q}\textbf{R}_{a}}\textbf{e}_{0}\hbox{\boldmath$\nabla$}_{r}V({\bf r}-{\bf R}_{a}).$ (22)
where $\omega_{ph}$ is the angular phonon frequency and q is its wave vector. The $a({\bf q})$ is the phonon annihilation operator (with corresponding momentum $\hbar\textbf{q}$ and energy $\hbar\omega_{ph}$), its adjoint is the phonon creation operator, $V({\bf r}-{\bf R}_{a})$ is the potential at the point ${\bf R_{a}}$, and ${\bf e}_{0}$ is the phonon polarization.
With above expressions, the Heisenberg equations (15) become a closed set of differential equations (“constitutive equations”) for $Y,C,D$, which can be obtained in explicit form using the following procedure:  [32, 33]
(i) setting up the Heisenberg equations of motion for the pair operators,
(ii) applying anti-commutation rules for the Fermion operators $\hat{c}_{j}^{c{\dagger}},\hat{c}_{j}^{c},\hat{d}_{j}^{v{\dagger}}$ and $\hat{d}_{j}^{v}$ to bring all operator products into normal order,
(iii) computing the expectation values of the relevant operators,
(iv) using an interpolation procedure to obtain a continuum dependence on the position variables (for example, Ref.[33]),
(v) making use of the RPA to factorize four-point density matrices.
As a result we obtain the constitutive equations for the interband transition density amplitudes, which for any couple of bands are of the following form
$-i\hbar\partial_{t}Y_{12}+H_{eh}Y_{12}+X^{Y}_{12}$ (23)
$={M_{0}}\left(E\delta_{12}-E_{1}C_{12}-E_{2}D_{21}\right)-i\hbar\left(\frac{\partial Y_{12}}{\partial t}\right)_{\rm irrev},$
with $X^{Y}_{12}$ the electron-hole exchange term, while the last dissipative term is responsible for irreversible excitation transfers. For intraband transitions (time dependence of the population of the band states) we obtain
$-i\hbar\partial_{t}C_{12}+H_{ee}C_{12}+X^{C}_{12}$ (24)
$=-{M_{0}}\left(E_{1}Y_{12}-E_{2}Y^{*}_{21}\right)-i\hbar\left(\frac{\partial C_{12}}{\partial t}\right)_{\rm irrev},$
$-i\hbar\partial_{t}D_{12}+H_{hh}D_{12}+X^{D}_{12}$ (25)
$=-{M_{0}}\left(Y_{21}E_{1}-Y^{*}_{12}E_{2}\right)-i\hbar\left(\frac{\partial D_{12}}{\partial t}\right)_{\rm irrev},$
where $X^{C}_{12}$ and $X^{D}_{12}$ are the electron-electron and hole-hole exchange terms, respectively. The numerical subscripts are abbreviations for the coordinates, such as in $Y_{12}=Y({\bf r}_{1},{\bf r}_{2})$ etc. The exchange terms $X^{Y},X^{C},X^{D}$ read
$X^{Y}_{12}=\frac{i}{\hbar}\int d^{3}r\,(V_{12}-V_{23})(Y_{13}C_{32}-D_{31}Y_{23}),$
$X^{C}_{12}=\frac{i}{\hbar}\int\,(V_{13}-V_{23})(Y^{*}_{31}Y_{32}+C_{13}C_{32}),$ (26)
$X^{D}_{12}=\frac{i}{\hbar}\int\,(V_{13}-V_{23})(Y^{*}_{31}Y_{32}+D_{13}D_{32}),$
where $V_{12}$ is the statistically screened Coulomb potential. [34] It can be seen that those terms represent 4-point correlations, but are not taken into account for 2-point correlation functions. It should be stressed that Hamiltonians are bilocal - they are defined in $({\bf r}_{1},{\bf r}_{2})$ space.
The electromagnetic fields ${\bf A}$, ${\bf E}$ and $\Phi$ which appear in (II.1)-(20) are self-consistent fields that include the induced contributions produced by the sources and those contained in $Y,C,D$. The spatial dependence of the functions $Y,C,D$ in the constitutive equations refers to a macroscopic scale. Microscopic structures can be taken into account by an appropriate choice of the parameters, as, for example, effective masses and $M_{0}$ in expressions (23-25). The above expressions must be solved simultaneously with the Maxwell field equations
$-c^{2}\epsilon_{0}\nabla\times\nabla\times{\bf E}-\epsilon_{o}\epsilon_{b}\ddot{\bf E}=\ddot{\bf P},$ (27)
where the polarization is given by expression (11), and $c$ is the speed of light. A theoretical scheme containing higher order correlations has been presented in Ref. [35], which also discusses phonon-assisted transitions.
As mentioned above, the term $H_{ph}$ in the Hamiltonian (17) describes the electron-crystal lattice interaction. Using perturbation calculus, we obtain the matrix elements of $H_{eh}$, under the condition of momentum conservation, and that of $H_{ph}$, which involves transfer of a specific momentum q. The calculated transition probability per unit time of a process in which the valence electron is scattered to the conduction state $\psi_{ck_{2}}$, and a photon of energy $\hbar\omega$ with a phonon of momentum $\textbf{q}=\textbf{k}_{1}-\textbf{k}_{2}$ and energy $\hbar\omega_{\textbf{q}}$ are both absorbed, gives the absorption coefficient in the case of the phonon absorption [36]
$\alpha_{ph,abs}(\omega)$ (28)
=$=\left\{\begin{array}[]{ccc}0,&\hbox{for}&\hbar\omega<E_{g}-\hbar\omega_{ph}\\ C_{1}(\hbar\omega-E_{g}+\hbar\omega_{ph})^{2}n_{\hbox{\scriptsize{q}}},&\hbox{for}&\hbar\omega>E_{g}+\hbar\omega_{ph}\end{array}\right.,$ (31)
where $n_{\hbox{\scriptsize{q}}}$ is the phonon occupation number representing the number of available phonons with wave vector q, and $C_{1}$ is a constant. By considering all values of q corresponding to energy $\hbar\omega_{ph}$, we use the distribution
$n_{ph}(\omega_{ph})=\frac{1}{\exp\frac{\hbar\omega_{ph}}{k_{B}{\mathcal{T}}}-1},$ (32)
where $k_{B}$ is the Boltzmann constant, and $\mathcal{T}$ is the temperature. The absorption coefficient in the case of phonon emission is given by
$\alpha_{ph,em}(\omega)$ (33)
=$=\left\{\begin{array}[]{cc}0,&\hbar\omega<E_{g}+\hbar\omega_{ph}\\ C_{1}(\hbar\omega-E_{g}-\hbar\omega_{ph})^{2}n_{ph},&\hbar\omega>E_{g}+\hbar\omega_{ph}\end{array}\right.,$ (36)
The case of exciton formation in an indirect transition can be illustrated as follows. Let us consider the case of parabolic non-degenerate energy bands with a maximum of the valence band at $\textbf{k}=0$, and a minimum of the conduction band at $\textbf{k}=\textbf{q}_{0}$, with the dispersion
$E_{c}(\textbf{k}_{e})=\frac{\hbar^{2}(\textbf{k}_{e}-\textbf{q}_{0})^{2}}{2m^{*}_{e}}+E_{g},$
$E_{v}(\textbf{k}_{h})=-\frac{\hbar^{2}\textbf{k}_{h}^{2}}{2m^{*}_{h}}.$ (37)
For $k_{ex}=q_{0}$ we obtain possible exciton energies
$E_{ex}(q_{0})=E_{g}-|E_{n\ell mH}|,$ (38)
(see also equation (II.3)) for states below the energy gap, and a continuum for states above the indirect gap. The absorption coefficients due to indirect exciton transitions in a process in which a photon and a phonon are simultaneously absorbed, and for the lowest exciton state, it is given by
$\alpha_{ph,abs}(\omega)=0\quad\hbox{for}\quad\hbar\omega<E_{g}-E_{1}-\hbar\omega_{ph},$ (39)
$\alpha_{ph,abs}(\omega)=C_{2}(\hbar\omega-E_{g}+E_{1}+\hbar\omega_{ph})^{\frac{1}{2}}n_{ph}$
$\hbox{for}\;\hbar\omega>E_{g}-E_{1}-\hbar\omega_{ph}.$
where $C_{2}$ is a constant [36], and we have written $E_{1}=|E_{100H}|$. The analogous expression for the absorption coefficient under the emission of a phonon has the form
$\alpha_{ph,em}(\omega)=0\quad\hbox{for}\quad\hbar\omega<E_{g}-E_{1}+\hbar\omega_{ph},$ (40)
$\alpha_{ph,em}(\omega)=C_{2}(\hbar\omega-E_{g}+E_{1}-\hbar\omega_{ph})^{\frac{1}{2}}n_{ph}$
$\hbox{for}\;\hbar\omega>E_{g}-E_{1}+\hbar\omega_{ph}.$
The total contribution of phonons to the absorption is then given by
$\alpha_{ph,total}=\alpha_{ph,abs}(\omega)+\alpha_{ph,em}(\omega).$ (41)
It can be seen from the above expressions that the effect of phonons is relevant when considering continuum states.
As discussed in Ref. [37, 38], the phonon density of states contains two local maxima at 20 meV and 60 meV, with a weighted average of approximately $\hbar\omega_{ph}=40$ meV. This simplified approach of taking an estimate of the average phonon energy provides a good fit to experimental data. In the same manner, as mentioned above, an integration of $n_{\textbf{q}}$ over all of values of q yields an average phonon number $n_{ph}$ that can be used in equations (33)-(40).
II.2 Two-photon absorption
We next adapt the described RDMA approach to the case of two-photon absorption. We assume that static external fields are absent, thus neglecting the vector potential $A$ and the scalar potential $\Phi^{e/h}$ in the Hamiltonian expressions (II.1-20). In expressions (23-25), the electromagnetic field $\bf E$ now includes two frequencies $\omega_{a}$ and $\omega_{b}$, and is written as
${\bf E}=\textbf{E}_{0a}\exp(i\textbf{k}_{\textbf{a}}{\textbf{R}}-i\omega_{a}t)$ (42)
@@ II.2 Two-photon absorption
$|\textbf{k}_{j}|=\frac{\omega_{j}}{c}\sqrt{\epsilon(\omega_{j})}=n_{j}\frac{\omega_{j}}{c},\quad j=a,b,$ (43)
and $n_{j}$ are refractive indices at the frequencies $\omega_{j}$, R is the electron-hole pair center-of-mass coordinate
${\bf R}={\bf R}_{12}=\frac{m_{h}{\bf r}_{1}+m_{e}{\bf r}_{2}}{m_{h}+m_{e}}.$ (44)
The linear optical properties are calculated by solving the interband equation (23), supplemented by the corresponding Maxwell equation, where the polarization (11) acts as a source. For computing the nonlinear optical properties we use the entire set of constitutive equations (23-24). Although finding a general solution of the equations is challenging, in special situations a solution can be found. For example, if one assumes that the matrices $Y,C$ and $D$ can be expanded in powers of the electric field ${\bf E}$, an iterative procedure can be used.
@@
    In the energy range $\hbar\omega_{a}\;+\hbar\omega_{b}\,>2E_{g}$, we solve the equations (II.3), (82) and following equations for the matrices $C,D$ assuming $V_{12}=0$ [24], thus entering the energy range represented by continuum states. The solution is obtained in terms of an appropriate Green function.
II.3 Discrete states - linear susceptibility
Our goal is to derive expressions for the NTA absorption coefficients. These can be obtained from the third-order nonlinear susceptibility, which, in turn, can be determined via an iterative procedure within the context of the RDMA. The first step in the iteration consists of solving equation (23), which at this stage takes on the form